### v1

@@ 6.1. Non-concentration inequalities
Recall $\mathfrak{g}=\mathfrak{sl}_{d+1}$. Write $\mathfrak{b}$ the Lie algebra of $B$, i.e. the subspace of upper triangular matrices in $\mathfrak{g}$. Note that $W$ is $\operatorname{Ad}(B)$-invariant if and only if it is $\operatorname{ad}(\mathfrak{b})$-invariant. Observe the relation $E_{i,j},E_{k,l}]={\mathbbm{1}}_{j=k}E_{i,l}-{\mathbbm{1}}_{i=l}E_{k,j}$ for all $i,j,k,l\in\{1,\dots,d+1\}$. In particular, for $i<j$ and $k<l$, the matrix $E_{i,j},E_{k,l}]$ is either 0 or up to a sign an elementary matrix located either to the right or above $E_{i,j}$. This justifies the “if” direction in Lemma˜6.8.
We now assume $W$ to be $\operatorname{Ad}(B)$-invariant and establish the announced decomposition. By invariance under diagonal matrices, $W$ must be of the form $W=E\bigoplus\oplus_{(i,j)\in S}\mathbb{R}E_{i,j}$ where $E$ is a subspace of diagonal matrices, and $S$ does not intersect the diagonal. If $E\neq\{0\}$, then by $\operatorname{ad}(\mathfrak{b})$-invariance, $W$ must contain a line $\mathbb{R}E_{i,i+1}$ for some $i\in\{1,\dots,d\}$. But the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by such a line has dimension at least $d$, which is absurd because $\dim W\leq d$. Hence $E=\{0\}$. Noting that for every $i>j$, the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by $\mathbb{R}E_{i,j}$ intersects the diagonal subspace, we further deduce $S\subseteq\{(i,j)\,:\,1\leq i<j\leq d+1\}$. The final claim on $S$ follows from $\operatorname{ad}(\mathfrak{b})$-invariance and the bracket relation exhibited in the first paragraph. ∎
We are finally able to show Proposition˜6.7, thus completing the proofs of Propositions 6.4, 6.6.
Proof of Proposition˜6.7.
@@
$\mathfrak{g}_{-}^{\prime}=\mathbb{R}E_{d+1,2}\oplus\cdots\oplus\mathbb{R}E_{d+1,d}.$
@@
To begin with, we let $k=\dim\operatorname{Proj}_{R_{1}}(W)\geq 1$. We choose
@@
Decompose $\mathfrak{g}_{-}=V_{0}\oplus V_{1}$, where $V_{0},V_{1}$ are linear subspaces defined by
$V_{0}=\mathbb{R}E_{d+1,1}\oplus\mathbb{R}E_{d+1,2}\oplus\cdots\oplus\mathbb{R}E_{d+1,d+1-k};$
$V_{1}=\mathbb{R}E_{d+1,d+2-k}\oplus\cdots\oplus\mathbb{R}E_{d+1,d}.$
Using Lemma˜6.8, we also decompose $W=\operatorname{Proj}_{R_{1}}(W)\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}}(W)$. Then we have
(6.8) $\operatorname{Ad}(g_{0})\mathfrak{g}_{-}\oplus W=\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}(W)}\oplus V_{1}\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}(W)}.$
Here we used that $V_{1}$ and $\operatorname{Ad}(g_{0})V_{1}$ coincide modulo $W$. Let
@@
Hence, we can apply the induction hypothesis (more precisely its equivalent version from Proposition˜6.6) to the pair $(G^{\prime},W^{\prime})$ to obtain $g^{\prime}_{1},\cdots,g^{\prime}_{d}\in G^{\prime}$ such that
(6.9) $\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}^{\prime}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}^{\prime}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}^{\prime}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}.$
@@
$C=\mathbb{R}E_{2,1}\oplus\mathbb{R}E_{3,1}\oplus\cdots\oplus\mathbb{R}E_{d+1,1}.$
@@
(6.10) $\operatorname{Ad}(g_{1}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\operatorname{Ad}(g_{2}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\cdots\oplus\operatorname{Ad}(g_{d}^{\prime\prime})\mathbb{R}E_{d+1,1}=C.$
@@
(6.11) $\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}\oplus C.$
@@
(6.12) $\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}}(W)\oplus\mathfrak{g}^{\prime}\oplus C=\mathfrak{g}$
@@ 6.2. Linear multislicing
Our aim is to show that $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ allow for a supercritical multislicing estimate, see Proposition˜6.14. For that, we first present a submodular inequality for covering numbers (Lemma˜6.10) and use it to connect $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with the properties of the individual projectors $\pi_{F_{i}}$ (Lemma˜6.11). We then deduce that a random subspace $F$ whose law $\Xi$ has the property $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ must enjoy both supercritical and subcritical projection theorems (Lemmas 6.12, 6.13). Those estimates refine that of Bourgain [13] and He [26] which were established under the stronger non-concentration condition that $\Xi$ satisfies $\operatorname{(MNC)}$ with $q=1$. From there, we use [6] to combine our projection theorems into the multislicing estimate Proposition˜6.14
@@
(6.13) ${\mathcal{N}}_{\mathcal{P}}(A){\mathcal{N}}_{\mathcal{Q}}(A)\geq\frac{c^{2}}{4L^{3}}{\mathcal{N}}_{\mathcal{R}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$
In the case where $L=1$, the result is due to [6, Lemma 2.6]. We deduce the refinement presented in Lemma˜6.10. Such an upgrade is convenient to deal with situations where partitions ${\mathcal{P}},{\mathcal{Q}},{\mathcal{R}},{\mathcal{S}}$ do not exactly fit together.
@@
Now consider ${\mathcal{P}}_{0}={\mathcal{P}}\vee{\mathcal{S}}$, ${\mathcal{Q}}_{0}={\mathcal{Q}}\vee{\mathcal{S}}$ and ${\mathcal{R}}_{0}={\mathcal{P}}\vee{\mathcal{Q}}\vee{\mathcal{S}}$. By [6, Lemma 2.6] applied to ${\mathcal{P}}_{0}$, ${\mathcal{Q}}_{0}$, ${\mathcal{R}}_{0}$ and $\mathcal{S}$}, there is a subset $A^{\prime}\subseteq A$ such that ${\mathcal{N}}_{{\mathcal{R}}_{0}}(A^{\prime})\geq(1-c){\mathcal{N}}_{{\mathcal{R}}_{0}}(A)$ and
(6.14) ${\mathcal{N}}_{{\mathcal{P}}_{0}}(A){\mathcal{N}}_{{\mathcal{Q}}_{0}}(A)\geq\frac{c^{2}}{4}{\mathcal{N}}_{{\mathcal{R}}_{0}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$
@@
(6.15) $\lVert F_{1}\wedge\dots\wedge F_{q}\rVert\geq r.$
@@
(6.16) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(\pi_{\oplus_{i}F_{i}}A),$
@@
(6.17) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{||F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(A)^{q-1}{\mathcal{N}}_{\rho}(\pi_{||\oplus_{i}F_{i}}A^{\prime})$
@@
The first inequality (6.16) is a simple counting, see e.g. [26, Lemma 15]. We focus on proving (6.17).
@@
$\lVert(F_{1}\oplus\dots\oplus F_{i})\wedge F_{i+1}\rVert\geq\lVert F_{1}\wedge\dotsc\wedge F_{q}\rVert\geq r,$
@@
From $\lVert F_{1}\wedge F_{2}\rVert\geq r$ we know that ${\mathcal{R}}\overset{r^{-O_{D}(1)}}{\prec}{\mathcal{P}}\vee{\mathcal{Q}}$ and it is always true that ${\mathcal{P}}\vee{\mathcal{Q}}\overset{O_{D}(1)}{\prec}{\mathcal{R}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{P}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{Q}}$. Thus, the inequality (6.17) follows from Lemma˜6.10. ∎
@@
(6.18) ${\mathcal{N}}_{\rho}(A)\geq\rho^{-D\alpha+\varepsilon},$
@@
(6.19) $\sup_{v\in\mathbb{R}^{D}}{\mathcal{N}}_{\rho}(A\cap B^{\mathbb{R}^{D}}_{r}(v))\leq\rho^{-\varepsilon}r^{c}{\mathcal{N}}_{\rho}(A).$
@@
$\begin{split}{\mathcal{E}}:=\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{-\alpha k-\varepsilon}\}\end{split}$
@@
Let $(q,m)$ be the couple of integers playing a role in the assumption $\operatorname{(MNC)}^{\perp}$ for $\Xi$. If $q=1$, the result is known. It is indeed the higher rank version of Bourgain’s projection theorem [13], due to the second-named author [26]. We deduce from there the general case $q\geq 1$. Note that throughout the proof, we may assume $A$ to be $2\rho$-separated. We may also allow the upper bound on $\rho$ to depend [^4] Indeed, if we establish the lemma for a pair $(\varepsilon,\rho)$ then it is automatically valid for $(\varepsilon^{\prime},\rho)$ with $\varepsilon^{\prime}\in(0,\varepsilon)$, because when passing from $\varepsilon$ to $\varepsilon^{\prime}$, assumptions get stronger and the conclusion gets weaker. on $\varepsilon$ (not only $D,c$). We let $\varepsilon_{1},\varepsilon_{2}>0$ be parameters to specify below in terms of $D$ and $c$. We use the shorthand $\mathcal{G}:=\operatorname{Gr}(\mathbb{R}^{D},k)$.
@@
${\mathcal{E}}_{1}:=\{\,{\underline{F}}\in\mathcal{G}^{q}:\|F_{1}^{\perp}\wedge\dots\wedge F_{q}^{\perp}\|\leq\rho^{(\varepsilon+\varepsilon_{1})/c}\}\,\,\,\text{ satisfies}\,\,\,\Xi^{\otimes q}({\mathcal{E}}_{1})\leq\rho^{\varepsilon_{1}}.$
@@
Let ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$. Up to assuming $\varepsilon\leq\varepsilon_{1}$, Equation (6.17) implies that for every set $S\subseteq A$, there exists a subset $S^{\prime}\subseteq S$ such that $|S^{\prime}|\geq\rho^{O_{D,c}(\varepsilon_{1})}|S|$ and
(6.20) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}S)\geq\rho^{O_{D,c}(\varepsilon_{1})}|S|^{q-1}{\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}S^{\prime}).$
Taking $S$ to be a not too small subset of $A$, we use (6.20) to obtain an explicit lower bound on $\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}S)$, see (6.22). As a lower bound on $|S|^{q-1}$ comes directly from the assumption (6.18), we focus on ${\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}S^{\prime})$.
Let $\Upsilon$ denote the restriction of $\Xi^{\otimes q}$ to $\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$, renormalised into a probability measure. Note that the random Dq(Dk)D-q(D-k)-plane $(\cap_{i}F_{i})_{{\underline{F}}\sim\Upsilon}$ satisfies the non-concentration condition $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon-\varepsilon_{1}},c)$ and $q=1$. Therefore, provided $\varepsilon+\varepsilon_{1},\varepsilon_{2}\lll_{D,c}1$ and $\rho\lll_{D,c}1$, there is an event ${\mathcal{E}}_{2}\subseteq\mathcal{G}^{q}$ such that $\Upsilon({\mathcal{E}}_{2})\leq\rho^{\varepsilon_{2}}$, and for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{2}$, for all $A^{\prime}\subseteq A$ with $|A^{\prime}|\geq\rho^{\varepsilon_{2}}|A|$, we have
(6.21) ${\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}A^{\prime})\geq\rho^{-\alpha(D-q(D-k))-\varepsilon_{2}}.$
Combining (6.20), (6.18) and (6.21), we obtain that for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus({\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$, and $A^{\prime\prime}\subseteq A$ with $|A^{\prime\prime}|\geq\rho^{\varepsilon_{1}}|A|$, we have
(6.22) $\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A^{\prime\prime})\geq\rho^{-\alpha k-\frac{1}{2q}\varepsilon_{2}},$
provided $\varepsilon\leq\varepsilon_{1}\lll_{D,c}\varepsilon_{2}\lll_{D,c}1$.
To conclude the proof, we argue by contradiction, assuming $\Xi({\mathcal{E}})>\rho^{\varepsilon}$. For each $F\in{\mathcal{E}}$, let $A^{\prime}_{F}\subseteq A$ be such that $\lvert A^{\prime}_{F}\rvert\geq\rho^{\varepsilon}\lvert A\rvert$ and ${\mathcal{N}}_{\rho}(\pi_{F}A^{\prime}_{F})<\rho^{-\alpha k-\varepsilon}$. By a Fubini argument such as [26, Lemma 19], $q$ independent copies of $A^{\prime}_{F}$ are likely to intersect in a rather large subset:
@@
${\mathcal{E}}_{3}:=\bigl{\{}\,{\underline{F}}\,:\,\lvert A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}\rvert\leq\rho^{2q\varepsilon}|A|\,\bigr{\}}\,\,\,\text{ satisfies}\,\,\,\Xi^{\otimes q}({\mathcal{E}}_{3})\leq 1-\rho^{4q\varepsilon}.$
Note that $\cup_{i=1}^{3}{\mathcal{E}}_{i}$ has $\Xi^{\otimes q}$-measure bounded above by $\rho^{\varepsilon_{1}}+\rho^{\varepsilon_{2}}+1-\rho^{4q\varepsilon}$ which is strictly less than $1$ provided $\varepsilon\lll_{D}\varepsilon_{1}\leq\varepsilon_{2}$ and $\rho\lll_{\varepsilon}1$. In particular, we may consider ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus\cup_{i=1}^{3}{\mathcal{E}}_{i}$. Setting $A^{\prime\prime}=A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}$, we have $|A^{\prime\prime}|\geq\rho^{2q\varepsilon}|A|$ because ${\underline{F}}\notin{\mathcal{E}}_{3}$, while the inclusions $A^{\prime\prime}\subseteq A^{\prime}_{F_{i}}$ yield

### v2

@@ 7.1. Non-concentration inequalities
Recall $\mathfrak{g}=\mathfrak{sl}_{d+1}$. Write $\mathfrak{b}$ the Lie algebra of $B$, i.e., the subspace of upper triangular matrices in $\mathfrak{g}$. As $B$ is Zariski-connected, the subspace $W$ is $\operatorname{Ad}(B)$-invariant if and only if it is $\operatorname{ad}(\mathfrak{b})$-invariant. Observe the relation $E_{i,j},E_{k,l}]={\mathbbm{1}}_{j=k}E_{i,l}-{\mathbbm{1}}_{i=l}E_{k,j}$ for all $i,j,k,l\in\{1,\dots,d+1\}$. In particular, for $i<j$ and $k<l$, the matrix $E_{i,j},E_{k,l}]$ is either 0 or up to a sign an elementary matrix located either to the right or above $E_{i,j}$. This justifies the “if” direction in Lemma˜7.8.
We now assume $W$ to be $\operatorname{Ad}(B)$-invariant and establish the announced decomposition. By invariance under diagonal matrices, $W$ must be of the form $W=E\oplus\operatorname{span}\{\,E_{i,j}\,;\,(i,j)\in S\,\}$ where $E$ is a subspace of diagonal matrices, and $S$ does not intersect the diagonal. If $E\neq\{0\}$, then by $\operatorname{ad}(\mathfrak{b})$-invariance, $W$ must contain a line $\mathbb{R}E_{i,i+1}$ for some $i\in\{1,\dots,d\}$. But the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by such a line has dimension at least $d$, which is absurd because $\dim W\leq d$. Hence $E=\{0\}$. Noting that for every $i>j$, the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by $\mathbb{R}E_{i,j}$ intersects the diagonal subspace, we further deduce $S\subseteq\{\,(i,j)\,:\,1\leq i<j\leq d+1\,\}$. The final claim on $S$ follows from $\operatorname{ad}(\mathfrak{b})$-invariance and the bracket relation exhibited in the first paragraph. ∎
We are finally able to show Proposition˜7.7, thus completing the proofs of Propositions 7.4, 7.6.
Proof of Proposition˜7.7.
@@
$\mathfrak{g}_{-}^{\prime}=\operatorname{span}\left\{E_{d+1,2},\dotsc,E_{d+1,d}\right\}.$
@@
To begin with, choose
@@
Now consider $k:=\dim\operatorname{Proj}_{R_{1}}(W)$, which, by Lemma˜7.8, satisfies $1\leq k\leq d$. We then decompose $\mathfrak{g}_{-}=V_{0}\oplus V_{1}$, where
$V_{0}:=\operatorname{span}\left\{\,E_{d+1,j}\,:\,1\leq j\leq d+1-k\,\right\},$
$V_{1}:=\operatorname{span}\left\{\,E_{d+1,j}\,:\,d+1-k<j\leq d\,\right\},$
so that $\dim V_{0}=d-(k-1)$ and $\dim V_{1}=k-1$. Using Lemma˜7.8, we also decompose $W=\operatorname{Proj}_{R_{1}}(W)\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}}(W)$. Observe that $V_{1}$ and $\operatorname{Ad}(g_{0})V_{1}$ coincide modulo $W$, leading to
$\operatorname{Ad}(g_{0})\mathfrak{g}_{-}\oplus W=\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}(W)}\oplus V_{1}\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}(W)}.$ (7.8)
Let
@@
Hence, we can apply the induction hypothesis (more precisely its equivalent version from Proposition˜7.6) to the pair $(G^{\prime},W^{\prime})$ to obtain $g^{\prime}_{1},\cdots,g^{\prime}_{d}\in G^{\prime}$ such that
$\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}^{\prime}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}^{\prime}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}^{\prime}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}.$ (7.9)
@@
$C:=\operatorname{span}\{\,E_{i,1}\,:\,2\leq i\leq d+1\,\}.$
@@
$\operatorname{Ad}(g_{1}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\operatorname{Ad}(g_{2}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\cdots\oplus\operatorname{Ad}(g_{d}^{\prime\prime})\mathbb{R}E_{d+1,1}=C.$ (7.10)
@@
$\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}\oplus C.$ (7.11)
@@
$\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}}(W)\oplus\mathfrak{g}^{\prime}\oplus C=\mathfrak{g}$ (7.12)
@@ 7.2. Linear multislicing
Our aim is to show that $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ allow for a supercritical multislicing estimate, see Proposition˜7.14. For that, we first present a submodular inequality for covering numbers (Lemma˜7.10) and use it to connect $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with the properties of the individual projectors $\pi_{F_{i}}$ (Lemma˜7.11). We then deduce that a random subspace $F$ whose law $\Xi$ has the property $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ must enjoy both supercritical and subcritical projection theorems (Lemmas 7.12, 7.13). Those estimates refine that of Bourgain [13] and He [29] which were established under the stronger non-concentration condition that $\Xi$ satisfies $\operatorname{(MNC)}$ with $q=1$. From there, we use [4] to combine our projection theorems into the multislicing estimate Proposition˜7.14
@@
${\mathcal{N}}_{\mathcal{P}}(A){\mathcal{N}}_{\mathcal{Q}}(A)\geq\frac{c^{2}}{4L^{3}}{\mathcal{N}}_{\mathcal{R}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$ (7.13)
In the case where $L=1$, the result is exactly [4, Lemma 2.6]. We deduce the refinement presented in Lemma˜7.10. Such an upgrade is convenient to deal with situations where partitions ${\mathcal{P}},{\mathcal{Q}},{\mathcal{R}},{\mathcal{S}}$ do not exactly fit together.
@@
Now consider ${\mathcal{P}}_{0}={\mathcal{P}}\vee{\mathcal{S}}$, ${\mathcal{Q}}_{0}={\mathcal{Q}}\vee{\mathcal{S}}$ and ${\mathcal{R}}_{0}={\mathcal{P}}\vee{\mathcal{Q}}\vee{\mathcal{S}}$. By [4, Lemma 2.6] applied to ${\mathcal{P}}_{0}$, ${\mathcal{Q}}_{0}$, ${\mathcal{R}}_{0}$ and $\mathcal{S}$}, there is a subset $A^{\prime}\subseteq A$ such that ${\mathcal{N}}_{{\mathcal{R}}_{0}}(A^{\prime})\geq(1-c){\mathcal{N}}_{{\mathcal{R}}_{0}}(A)$ and
${\mathcal{N}}_{{\mathcal{P}}_{0}}(A){\mathcal{N}}_{{\mathcal{Q}}_{0}}(A)\geq\frac{c^{2}}{4}{\mathcal{N}}_{{\mathcal{R}}_{0}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$ (7.14)
@@
$\lVert F_{1}\wedge\dots\wedge F_{q}\rVert\geq r.$ (7.15)
@@
$\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(\pi_{\oplus_{i}F_{i}}A),$ (7.16)
@@
$\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{||F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(A)^{q-1}{\mathcal{N}}_{\rho}(\pi_{||\oplus_{i}F_{i}}A^{\prime})$ (7.17)
@@
The first inequality (7.16) is a simple counting, see e.g. [29, Lemma 15]. We focus on proving (7.17).
@@
$\lVert(F_{1}\oplus\dots\oplus F_{q-1})\wedge F_{q}\rVert\geq\lVert F_{1}\wedge\dotsc\wedge F_{q}\rVert\geq r,$
@@
From $\lVert F_{1}\wedge F_{2}\rVert\geq r$ we know that ${\mathcal{R}}\overset{r^{-O_{D}(1)}}{\prec}{\mathcal{P}}\vee{\mathcal{Q}}$ and it is always true that ${\mathcal{P}}\vee{\mathcal{Q}}\overset{O_{D}(1)}{\prec}{\mathcal{R}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{P}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{Q}}$. Thus, the inequality (7.17) follows from Lemma˜7.10. ∎
@@
${\mathcal{N}}_{\rho}(A)\geq\rho^{-D\alpha+\varepsilon},$ (7.18)
@@
$\sup_{v\in\mathbb{R}^{D}}{\mathcal{N}}_{\rho}(A\cap B^{\mathbb{R}^{D}}_{r}(v))\leq\rho^{-\varepsilon}r^{c}{\mathcal{N}}_{\rho}(A).$ (7.19)
@@
$\begin{split}{\mathcal{E}}:=\bigl\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{-\alpha k-\varepsilon}\,\bigr\}\end{split}$
@@
Let $(q,m)$ be the pair of integers playing a role in the assumption $\operatorname{(MNC)}^{\perp}$ for $\Xi$. If $q=1$, the result is known. It is indeed the higher rank version of Bourgain’s projection theorem [13], due to the second-named author [29]. We deduce from there the general case $q\geq 1$.
Note that throughout the proof, we may assume $A$ to be $2\rho$-separated, so that ${\mathcal{N}}_{\rho}(S)=\lvert S\rvert$ holds for any of its subsets $S\subseteq A$. We may also allow the upper bound on $\rho$ to depend [^5] Indeed, if we establish the lemma for a pair $(\varepsilon,\rho)$ then it is automatically valid for $(\varepsilon^{\prime},\rho)$ with $\varepsilon^{\prime}\in(0,\varepsilon)$, because when passing from $\varepsilon$ to $\varepsilon^{\prime}$, assumptions get stronger and the conclusion gets weaker. on $\varepsilon$ (not only $D,c$). We let $\varepsilon_{1},\varepsilon_{2}>0$ be parameters to specify below in terms of $D$ and $c$. We use the shorthand $\mathcal{G}:=\operatorname{Gr}(\mathbb{R}^{D},k)$.
@@
${\mathcal{E}}_{1}:=\bigl\{\,{\underline{F}}\in\mathcal{G}^{q}:\|F_{1}^{\perp}\wedge\dots\wedge F_{q}^{\perp}\|\leq\rho^{(\varepsilon+\varepsilon_{1})/c}\,\bigr\}\text{ satisfies }\Xi^{\otimes q}({\mathcal{E}}_{1})\leq\rho^{\varepsilon_{1}}.$
@@
Let ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$. Up to assuming $\varepsilon\leq\varepsilon_{1}$, Equation (7.17) applied to the family $(F_{i}^{\perp})_{i=1}^{q}$ implies that for every set $S\subseteq A$, there exists a subset $S^{\prime}\subseteq S$ such that $\lvert S^{\prime}\rvert\geq\rho^{O_{D,c}(\varepsilon_{1})}\lvert S\rvert$ and
@@
$\prod_{i=1}^{q}{\mathcal{N}}_{\rho}\bigl(\pi_{F_{i}}S\bigr)\geq\rho^{O_{D,c}(\varepsilon_{1})}\lvert S\rvert^{q-1}{\mathcal{N}}_{\rho}(\pi_{\bigcap_{i}F_{i}}S^{\prime}),$
which in particular yields